# supplement.pdf

### v2

@@ Abstract
Keywords: Mediation analysis; High-dimensional inference; Bias correction; Composite hypothesis; Super-efficiency.
@@ 1 Introduction
Originally motivated in psychology (Baron and Kenny, 1986), mediation analysis has found widespread application in various scientific fields, including medicine, genomics, economics, and many others (Rijnhart et al., 2021; Zeng et al., 2021; Celli, 2022; Preacher, 2015), over the past few decades. Given independent and identically distributed (i.i.d.) triples $(Y_{i},A_{i},M_{i})_{i=1}^{n}$, mediation analysis focuses on examining the effect of exposures $A_{i}\in\mathbb{R}^{q}$ on an outcome $Y_{i}\in\mathbb{R}$, which may be mediated by some potential intermediate variables $M_{i}\in\mathbb{R}^{p}$ known as mediators. In the case where the relationships among $Y_{i}$, $A_{i}$, and $M_{i}$ are linear, one can consider the following linear structural equation model (LSEM):
@@
for each $i=1,\ldots,n$. Here, both $E_{i}\in\mathbb{R}^{p}$ and $Z_{i}\in\mathbb{R}$ are random noise. The matrix $\beta_{A}\in\mathbb{R}^{p\times q}$ and the vectors $\theta_{A}\in\mathbb{R}^{q}$ and $\theta_{M}\in\mathbb{R}^{p}$ contain the unknown regression coefficients that encode the relationships among the exposures, mediators and outcome. In this paper, we consider the high-dimensional setting where the dimension $p$ of the mediators diverges and the dimension $q$ of the exposures may also grow with the sample size $n$. Assuming no unmeasured confounders, the joint mediated effect through the mediators, also referred to as the natural indirect effect or mediation effect, is captured by the parameter $\gamma=\beta_{A}^{\top}\theta_{M}\in\mathbb{R}^{q}$ (Robins and Greenland, 1992; Pearl, 2001; VanderWeele and Vansteelandt, 2014). Our focus is on the hypothesis testing problem,
@@
where $\textbf{0}_{q}\in\mathbb{R}^{q}$ is the vector of all zeros. Due to its practical importance, numerous statistical methods have been proposed for the problem (3), primarily in the context of low-dimensional mediators (MacKinnon et al., 2002; VanderWeele and Vansteelandt, 2009, 2014). For comprehensive reviews, we refer to the survey papers Ten Have and Joffe (2012); Preacher (2015); Barfield et al. (2017) and the monograph VanderWeele (2015). Nowadays, high-dimensional data are ubiquitous in many areas, such as bioinformatics (Johnstone and Titterington, 2009; Schadt et al., 2010; Zeng et al., 2021). This has led to a growing need for new statistical methods for mediation analysis with high-dimensional mediators, where the number of potential mediators may be comparable to, or even larger than, the sample size. For instance, genome-wide association studies have investigated the impact of early-life trauma on cortisol stress reactivity in adulthood through hundreds of thousands of DNA methylation levels (Houtepen et al., 2016; Guo et al., 2022a). Epidemiological studies have also confirmed the role of socioeconomic factors, mediated through molecular-level traits including methylation, in disease susceptibility (Tobi et al., 2018; Huang, 2019). In neuroscience, there is interest in identifying brain regions, comprising a large volume of voxels, whose activity levels act as potential mediators in the relationship between a thermal stimulus and self-reported pain (Chén et al., 2018). At first glance, in mediation analysis, one might attempt to address $\beta_{A}$ and $\theta_{M}$ separately through their respective equations in the LSEM. However, the coexistence of $\beta_{A}\neq 0$ and $\theta_{M}\neq 0$ does not imply $\gamma\neq 0$. In addition, separately analyzing the two equations do not account for the complex interplay between the exposure, mediator and outcome. Consequently, existing techniques for high-dimensional linear models (Zhang and Zhang, 2014; Ning and Liu, 2017; Zhu and Bradic, 2018; Javanmard and Lee, 2020, among many others) cannot be readily adapted to analyze the two coupled equations in the LSEM. Various methodologies have been proposed in response to the unique challenge of mediation analysis, especially in the presence of high-dimensional mediators. For example, one might first reduce the number of mediators through dimension reduction techniques, such as principal component analysis (Huang and Pan, 2016; Chén et al., 2018) and variable screening (Zhang et al., 2016). Alternatively, studies such as Huang (2019); Dai et al. (2022); Liu et al. (2022) consider testing whether $\beta_{A,j}\theta_{M,j}=0$ for all $j=1,\ldots,p$, where $\beta_{A,j}$ and $\theta_{M,j}$ are respectively the $j$th coordinates of $\beta_{A}$ and $\theta_{M}$. However, while these individual null hypotheses are meaningful in their own right, they do not collectively equate to $\gamma=0$. In addition, Huang (2019) requires a sparse $\beta_{A}$, whereas Dai et al. (2022) and Liu et al. (2022) impose implicit conditions on $\beta_{A}$ through an assumption of the weak dependence across coordinates $1,\ldots,p$. In contrast, The recent works by Zhou et al. (2020) and Guo et al. (2022b) directly address the overall mediation effect $\gamma$. Specifically, Zhou et al. (2020) proposed a debiased estimator and a test for the mediation effect, under some structural assumptions on the exposure–mediator coefficients $\beta_{A}$ to ensure consistency of the estimator. Guo et al. (2022b), observing that the mediation effect $\gamma$ is the difference between the total effect $\gamma+\theta_{A}$ and the natural direct effect $\theta_{A}$, proposed to estimate $\gamma$ by the difference between an estimator of the total effect and an estimator of the natural direct effect, and developed a Wald test for the mediation effect. Their method relies on the sign consistency of the estimated mediator–outcome coefficient $\theta_{M}$, which demands relatively strong assumptions, such as the uniform signal strength condition (Zhang and Zhang, 2014) and the irrepresentable condition (Zhao and Yu, 2006). As a consequence, the exposure–mediator coefficient $\beta_{A}$ in Guo et al. (2022b) cannot be too large in magnitude. The null hypothesis $\gamma=0$ includes the case that both $\beta_{A}=0$ and $\theta_{M}=0$, which is not that uncommon in applications such as genome-wide studies due to extreme sparse signals (Huang, 2019). In addition to structural assumptions on $\beta_{A}$, both tests of Zhou et al. (2020) and Guo et al. (2022b) do not address this peculiar case of the null hypothesis. The fundamental cause is that, when $\beta_{A}=0$ and $\theta_{M}=0$, the standard deviations of their estimators for the mediation effect decay to zero at a rate faster than $n^{-1/2}$, rendering their asymptotic normality results invalid; see a numerical demonstration in Section 4. In fact, the case of $\beta_{A}=0$ and $\theta_{M}=0$ is nontrivial even in the low-dimensional setting (Sobel, 1982; Barfield et al., 2017; Zhou et al., 2020). As a major contribution of this paper, we develop a test for the overall mediation effect $\gamma=0$ in the presence of high-dimensional mediators, with the following distinctive features. First, it remains valid in all cases of the null hypothesis, even in the challenging case of $\beta_{A}=0$ and $\theta_{M}=0$. Second, it does not impose assumptions on $\beta_{A}$, and thus is able to accommodate arbitrary exposure–mediator coefficients $\beta_{A}$, including a dense vector $\beta_{A}$. See Table 1 for a contrast between our method and the others. To the best of our knowledge, our test is the first of its kind to enjoy both of these features in high-dimensional mediation analysis. Our test is built on a novel debiased estimator for the mediation effect $\gamma$, achieved by adapting the technique of the variance-enhancement projection direction (Cai et al., 2021). Although the method of Zhou et al. (2020) is also based on debiasing a pilot estimator by projecting the sum of residuals, the projection direction in Zhou et al. (2020) is constructed to solely alleviate the bias of the pilot estimator. In contrast, in our procedure, in addition to correcting the bias, we construct the direction to also control the variance of the estimator. Therefore, the variance of our estimator would dominate the corresponding bias for any $\beta_{A}$, thus allowing for arbitrary exposure–mediation associations; see the discussion right after (12) for details. Compared to the work presented in Cai et al. (2021), which primarily concentrates on testing a linear contrast with a predetermined high-dimensional loading vector, the development of our test method confronts a distinctive theoretical hurdle. First, this challenge stems from the consideration of a random high-dimensional loading vector, which emerges due to the estimation of the unknown coefficient vector $\beta_{A}$. Second, the intricate interdependence between the two high-dimensional equations within the LSEM introduces a significant complication in our theoretical analysis. This complexity is distinct from the typical situation encountered in the high-dimensional linear models, where a single equation is usually addressed (e.g., Cai et al., 2021).
@@
Table 1: Validity of our method and the competing methods, and requirement on sign consistency.
@@
The rest of the paper is organized as follows. In Section 2, we present the proposed debiased estimator and the corresponding test for mediation effect. Theoretical investigations of the proposed test are provided in Section 3. Numerical studies on the proposed method and comparisons with other methods are presented in Section 4. In Section 5, we showcase the proposed method in a real data application, which investigates the mediation effect between smoking status, DNA methylation, and lung cancer development. The paper concludes with a final remark in Section 6. Notation. For a random variable/vector $X$, $\mathbb{E}[X]$ and $\mathrm{Var}[X]$ denote the expectation and the variance/covariance matrix of $X$, respectively. For a vector $x$, $x^{\top}$ is its transpose, $\|x\|_{r}$ represents its $\ell_{r}$ norm with $r=1,2,\infty$, and $\text{diag}(x)$ denotes the diagonal matrix whose diagonal is $x$. For a matrix $X$, $\|X\|_{2}$ and $\|X\|_{\infty}$ represent its spectral norm and element-wise $\ell_{\infty}$ norm, respectively. Let $X_{i}$ (respectively, $X_{\cdot j}$) be its $i$th row (respectively, $j$th column) and $X_{jk}$ represents its $(j,k)$th element. In particular, for the matrix $\beta_{A}$ in (1), $\beta_{A,j}$ is its $j$th column and $\beta_{A,jk}$ is its $(j,k)$th element. If $X$ is a squared matrix, $\Lambda_{\min}(X)$ and $\Lambda_{\max}(X)$ denote its smallest and largest eigenvalues, respectively. $\textbf{1}_{q}$ and $\textbf{0}_{q}$ are the vectors consisting of all ones and all zeros, respectively, in $\mathbb{R}^{q}$. Occasionally, we may omit the subscript $q$ when it’s clear from the context. $I_{n}$ represents the identity matrix of size $n$. $\mathbb{I}_{\left\{\cdot\right\}}$ denotes the indicator function. $N_{q}(\mu,\Sigma)$ represents the $q$-dimensional Gaussian distribution with mean $\mu$ and covariance matrix $\Sigma$, while $N(0,1)$ represents the standard Gaussian distribution in $\mathbb{R}$. For two non-negative sequences $\{a_{n}\}$ and $\{b_{n}\}$, we write $a_{n}\lesssim b_{n}$ (respectively, $a_{n}\gtrsim b_{n}$) if there is a constant $c>0$ not depending on $n$, such that $a_{n}\leq cb_{n}$ (respectively, $a_{n}\geq cb_{n}$) for all sufficiently large $n$. We write $a_{n}\asymp b_{n}$ if and only if both $a_{n}\lesssim b_{n}$ and $a_{n}\gtrsim b_{n}$. Moreover, $a_{n}\ll b_{n}$ if $a_{n}/b_{n}\to 0$ as $n\to\infty$.
@@ 2.1 Estimating Mediation Effect via a Debiased Approach
and $E_{M,i}=E_{i}^{\top}\theta_{M}$ for $i=1,\ldots,n$. With high-dimensional mediators, the first term $\hat{\Sigma}_{A}^{-1}\hat{\Sigma}_{AM}(\hat{\theta}_{M}-\theta_{M})$ in (7) becomes a non-negligible bias of $\tilde{\gamma}$ due to the high-dimensional penalty. To eliminate this bias, we employ a bias correction approach, as follows. Let $\tilde{g}_{j}$ be the $j$th column of $(\hat{\Sigma}_{A}^{-1}\hat{\Sigma}_{AM})^{\top}\in\mathbb{R}^{p\times q}$ and thus $\hat{\Sigma}_{A}^{-1}\hat{\Sigma}_{AM}(\hat{\theta}_{M}-\theta_{M})=\big{(}\tilde{g}_{1}^{\top}(\hat{\theta}_{M}-\theta_{M}),\ldots,\tilde{g}_{q}^{\top}(\hat{\theta}_{M}-\theta_{M})\big{)}^{\top}$. To estimate each coordinate of the bias, we first observe that, for any projection direction $u\in\mathbb{R}^{q+p}$,
@@
where the upper bound may be minimized by choosing an appropriate initial estimator $\hat{\theta}$ and imposing constraints on $\|\hat{\Sigma}_{X}u-g_{j}\|_{\infty}$ as in (9). For the second term $n^{-1}u^{\top}X^{\top}Z$ of (8), as its mean is zero, we find an effective $u$ to minimize its variance which is in the order of $u^{\top}\hat{\Sigma}_{X}u$. In addition, to facilitate statistical inference, an effective projection direction $u$ should ensure that the variance of the associated debiased estimator asymptotically dominates its bias, for all realizations of $g_{j}$. All of these considerations inspire us to adapt the variance-enhancement projection direction technique (Cai et al., 2021). Specifically, for each $j=1,\ldots,q$, the projection direction $\hat{u}_{j}$ is
@@
where $\lambda\asymp\sqrt{\log(q+p)/n}$ and $\mu\asymp\log n$ are two tuning parameters. Details on solving this constrained optimization and selecting the tuning parameters can be found in Section 4.1 of Guo et al. (2021a). Consequently, with $\hat{U}=(\hat{u}_{1},\ldots,\hat{u}_{q})\in\mathbb{R}^{(p+q)\times q}$, the estimated bias term in (7) is $n^{-1}\hat{U}^{\top}X^{\top}(X\hat{\theta}-Y)$, based on which we propose the following debiased estimator for $\gamma$,
@@
In the above, the constraint (9) is introduced to tackle the first term on the right-hand side of (8); this term may be viewed as the further bias associated with the estimator $n^{-1}u^{\top}X^{\top}(X\hat{\theta}-Y)$ for the $j$th coordinate $\tilde{g}_{j}^{\top}(\hat{\theta}_{M}-\theta_{M})$ of the bias in (7). This idea of minimizing the variance of a projected sum of errors while constraining the bias is a common practice in high-dimensional statistical inference with bias correction (Zhang and Zhang, 2014; Javanmard and Montanari, 2014). Cai et al. (2021) further introduced an additional constraint (10) to ensure that the variance of the second term on the right-hand side of (8) asymptotically dominates the first term for a fixed counterpart of $g_{j}$, and hence the variance of the debiased estimator would asymptotically dominate its bias. In our scenario, each $g_{j}$ is random due to estimating the parameter $\beta_{A}$ of the additional equation (4). This distinction, combined with the presence of two coupled equations in our model, sets our theoretical analysis apart from previous works. The constraint (11) is primarily a technical requirement to ensure that the error terms in (8) satisfy the Lindeberg’s condition.

### v3

@@ Abstract
keywords:
High-dimensional inference, Bias correction, Mediation analysis, Composite hypothesis, Super-efficiency
@@ 1 Introduction
Originally motivated in psychology [1], mediation analysis has found widespread application in various scientific fields, including medicine, genomics, economics, and many others [2, 3], over the past few decades. Given independent and identically distributed (i.i.d.) triples $(Y_{i},A_{i},M_{i})_{i=1}^{n}$, mediation analysis focuses on examining the effect of exposures $A_{i}\in\mathbb{R}^{q}$ on an outcome $Y_{i}\in\mathbb{R}$, which may be mediated by some potential intermediate variables $M_{i}\in\mathbb{R}^{p}$ known as mediators. In the case where the relationships among $Y_{i}$, $A_{i}$, and $M_{i}$ are linear, one can consider the following linear structural equation model (LSEM):
@@
for each $i=1,\ldots,n$. Here, both $E_{i}\in\mathbb{R}^{p}$ and $Z_{i}\in\mathbb{R}$ are random noise. The matrix $\beta_{A}\in\mathbb{R}^{p\times q}$ and the vectors $\theta_{A}\in\mathbb{R}^{q}$ and $\theta_{M}\in\mathbb{R}^{p}$ contain the unknown regression coefficients that encode the relationships among the exposures, mediators and outcome. In this paper, we consider the high-dimensional setting where the dimension $p$ of the mediators may diverge with the sample size $n$.
In the LSEM (1) and (2), the joint mediated effect through the mediators, also referred to as the natural indirect effect or mediation effect, is captured by the parameter $\gamma=\beta_{A}^{\top}\theta_{M}\in\mathbb{R}^{q}$ [4, 5]. It is known that, if the LSEM (1) and (2) is correctly specified with the absence of measured baseline covariates, $\gamma$ admits a causal interpretation under a counterfactual framework; see Appendix A for details. Our focus is on the hypothesis testing problem,
@@
where $\textbf{0}_{q}\in\mathbb{R}^{q}$ is the vector of all zeros. Due to its practical importance, numerous statistical methods have been proposed for the problem (3), primarily in the context of low-dimensional mediators [6, 5].
Nowadays, high-dimensional data are ubiquitous in many areas, such as bioinformatics [2]. This has led to a growing need for new statistical methods for mediation analysis with high-dimensional mediators, where the number of potential mediators may be comparable to, or even larger than, the sample size. For instance, genome-wide association studies have investigated the impact of early-life trauma on cortisol stress reactivity in adulthood through hundreds of thousands of DNA methylation levels [7, 8]. Epidemiological studies have also confirmed the role of socioeconomic factors, mediated through molecular-level traits including methylation, in disease susceptibility [9, 10]. In neuroscience, there is interest in identifying brain regions, comprising a large volume of voxels, whose activity levels act as potential mediators in the relationship between a thermal stimulus and self-reported pain [11]. At first glance, in mediation analysis, one might attempt to address $\beta_{A}$ and $\theta_{M}$ separately through their respective equations in the LSEM. However, the coexistence of $\beta_{A}\neq 0$ and $\theta_{M}\neq 0$ does not imply $\gamma\neq 0$. In addition, separately analyzing the two equations do not account for the complex interplay between the exposure, mediator and outcome. Consequently, existing techniques for high-dimensional linear models cannot be readily adapted to analyze the two coupled equations in the LSEM.
Various methodologies have been proposed in response to the unique challenge of mediation analysis, especially in the presence of high-dimensional mediators. For example, one might first reduce the number of mediators through dimension reduction techniques, such as principal component analysis [12, 11] and variable screening [13]. Alternatively, studies such as [10, 14, 15] consider testing whether $\beta_{A,j}\theta_{M,j}=0$ for all $j=1,\ldots,p$, where $\beta_{A,j}$ and $\theta_{M,j}$ are respectively the $j$th coordinates of $\beta_{A}$ and $\theta_{M}$. However, while these individual null hypotheses are meaningful in their own right, they do not collectively equate to $\gamma=0$. In addition, [10] requires a sparse $\beta_{A}$, whereas [14] and [15] impose implicit conditions on $\beta_{A}$ through an assumption of the weak dependence across coordinates $1,\ldots,p$. In contrast, The recent works by [16] and [17] directly address the overall mediation effect $\gamma$. Specifically, [16] proposed a debiased estimator and a test for the mediation effect, under some structural assumptions on the exposure–mediator coefficients $\beta_{A}$ to ensure consistency of the estimator. [17], observing that the mediation effect $\gamma$ is the difference between the total effect $\gamma+\theta_{A}$ and the natural direct effect $\theta_{A}$, proposed to estimate $\gamma$ by the difference between an estimator of the total effect and an estimator of the natural direct effect, and developed a Wald test for the mediation effect. Their method relies on the sign consistency of the estimated mediator–outcome coefficient $\theta_{M}$, which demands relatively strong assumptions, such as the uniform signal strength condition [18] and the irrepresentable condition [19]. As a consequence, the exposure–mediator coefficient $\beta_{A}$ in [17] cannot be too large in magnitude.
The null hypothesis $\gamma=0$ includes the case that both $\beta_{A}=0$ and $\theta_{M}=0$, which is not that uncommon in applications such as genome-wide studies due to extreme sparse signals [10, 21]. In addition to structural assumptions on $\beta_{A}$, both tests of [16] and [17] do not address this peculiar case of the null hypothesis. The fundamental cause is that, when $\beta_{A}=0$ and $\theta_{M}=0$, the standard deviations of their estimators for the mediation effect decay to zero at a rate faster than $n^{-1/2}$, rendering their asymptotic normality results invalid; see a numerical demonstration in Section 4. In fact, the case of $\beta_{A}=0$ and $\theta_{M}=0$ is nontrivial even in the low-dimensional setting [20, 21, 16].
As a major contribution of this paper, we develop a test for the overall mediation effect $\gamma=0$ in the presence of high-dimensional mediators, with the following distinctive features. First, it does not impose assumptions on $\beta_{A}$, and thus is able to accommodate arbitrary exposure–mediator coefficients $\beta_{A}$, including a dense vector $\beta_{A}$. Second, it remains valid in all cases of the null hypothesis, even in the challenging case of $\beta_{A}=0$ and $\theta_{M}=0$. See Table 1 for a contrast between our method and the others. To the best of our knowledge, our test is the first of its kind to enjoy both of these features in high-dimensional mediation analysis.
Our test is built on a novel debiased estimator for the mediation effect $\gamma$, achieved by adapting the technique of the variance-enhancement projection direction (VePD) [22]. Although the method of [16] is also based on debiasing a pilot estimator by projecting the sum of residuals, the projection direction in [16] is constructed to solely alleviate the bias of the pilot estimator. In contrast, in our procedure, in addition to correcting the bias, we construct the direction to also control the variance of the estimator. Therefore, the variance of our estimator would dominate the corresponding bias for any $\beta_{A}$, thus allowing for arbitrary exposure–mediation associations; see the discussion right after (12) for details. Compared to the work presented in [22], which primarily concentrates on testing a linear contrast with a predetermined high-dimensional loading vector, the development of our test method confronts a distinctive theoretical hurdle. First, this challenge stems from the consideration of a random high-dimensional loading vector, which emerges due to the estimation of the unknown coefficient vector $\beta_{A}$. Second, the intricate interdependence between the two high-dimensional equations within the LSEM introduces a significant complication in our theoretical analysis. This complexity is distinct from the typical situation encountered in the high-dimensional linear models, where a single equation is usually addressed [e.g., 22].
@@
Table 1: Validity of our method and the competing methods, and requirement on sign consistency
@@
00footnotetext: Note: While there is no direct sparsity condition on $\beta_{A}$ in [16], Assumption 2 therein imposes a structural requirement on the covariance structures of the mediators and exposures. The discussion following that assumption mentions that such requirement is related to the irrepresentable condition of [19]. When $\beta_{A}$ is dense, exposures and mediators are likely to be strongly correlated, making the irrepresentable condition hard to meet [19].
The rest of the paper is organized as follows. In Section 2, we present the proposed debiased estimator and the corresponding test for mediation effect. Theoretical investigations of the proposed test are provided in Section 3. Numerical studies on the proposed method and comparisons with other methods are presented in Section 4. In Section 5, we showcase the proposed method in a real data application, which investigates the mediation effect between smoking status, DNA methylation, and lung cancer development. The paper concludes with a final remark in Section 6.
Notation. For a vector $x$, $x^{\top}$ is its transpose, $\|x\|_{r}$ represents its $\ell_{r}$ norm with $r=1,2,\infty$, and $\text{diag}(x)$ denotes the diagonal matrix whose diagonal is $x$. For a matrix $X$, $\|X\|_{2}$ and $\|X\|_{\infty}$ represent its spectral norm and element-wise $\ell_{\infty}$ norm, respectively. Let $X_{i}$ (respectively, $X_{\cdot j}$) be its $i$th row (respectively, $j$th column) and $X_{jk}$ represents its $(j,k)$th element. In particular, for the matrix $\beta_{A}$ in (1), $\beta_{A,j}$ is its $j$th column and $\beta_{A,jk}$ is its $(j,k)$th element. If $X$ is a squared matrix, $\Lambda_{\min}(X)$ and $\Lambda_{\max}(X)$ denote its smallest and largest eigenvalues, respectively. $\textbf{1}_{q}$ and $\textbf{0}_{q}$ are the vectors consisting of all ones and all zeros, respectively, in $\mathbb{R}^{q}$. Occasionally, we may omit the subscript $q$ when it’s clear from the context. $I_{n}$ represents the identity matrix of size $n$. $\mathbb{I}_{\left\{\cdot\right\}}$ denotes the indicator function. $N_{q}(\mu,\Sigma)$ represents the $q$-dimensional Gaussian distribution with mean $\mu$ and covariance matrix $\Sigma$, while $N(0,1)$ represents the standard Gaussian distribution in $\mathbb{R}$. For two non-negative sequences $\{a_{n}\}$ and $\{b_{n}\}$, we write $a_{n}\lesssim b_{n}$ (respectively, $a_{n}\gtrsim b_{n}$) if there is a constant $c>0$ not depending on $n$, such that $a_{n}\leq cb_{n}$ (respectively, $a_{n}\geq cb_{n}$) for all sufficiently large $n$. We write $a_{n}\asymp b_{n}$ if and only if both $a_{n}\lesssim b_{n}$ and $a_{n}\gtrsim b_{n}$. Moreover, $a_{n}\ll b_{n}$ if $a_{n}/b_{n}\to 0$ as $n\to\infty$.
@@ 2.1 Estimating Mediation Effect via a Debiased Approach
and $E_{M,i}=E_{i}^{\top}\theta_{M}$ for $i=1,\ldots,n$. Intuitively, the second term on the right-hand side of (7) is asymptotically normal under regularity conditions. In contrast, with high-dimensional mediators, the first term $\hat{\Sigma}_{A}^{-1}\hat{\Sigma}_{AM}(\hat{\theta}_{M}-\theta_{M})$ in (7) becomes a non-negligible bias of $\tilde{\gamma}$ due to the high-dimensional penalty. To eliminate this bias, we employ a bias correction approach, as follows.
Let $\tilde{g}_{j}$ be the $j$th column of $(\hat{\Sigma}_{A}^{-1}\hat{\Sigma}_{AM})^{\top}\in\mathbb{R}^{p\times q}$ and thus $\hat{\Sigma}_{A}^{-1}\hat{\Sigma}_{AM}(\hat{\theta}_{M}-\theta_{M})=\big{(}\tilde{g}_{1}^{\top}(\hat{\theta}_{M}-\theta_{M}),\ldots,\tilde{g}_{q}^{\top}(\hat{\theta}_{M}-\theta_{M})\big{)}^{\top}$. To estimate each coordinate of the bias, we first observe that, for any projection direction $u\in\mathbb{R}^{q+p}$,
@@
where the upper bound may be minimized by choosing an appropriate initial estimator $\hat{\theta}$ and imposing constraints on $\|\hat{\Sigma}_{X}u-g_{j}\|_{\infty}$ as in (9). For the second term $n^{-1}u^{\top}X^{\top}Z$ of (8), as its mean is zero, we find an effective $u$ to minimize its variance which is in the order of $u^{\top}\hat{\Sigma}_{X}u$. In addition, to facilitate statistical inference, an effective projection direction $u$ should ensure that the variance of the associated debiased estimator asymptotically dominates its bias, for all realizations of $g_{j}$. All of these considerations inspire us to adapt the VePD technique [22].
Specifically, for each $j=1,\ldots,q$, the projection direction $\hat{u}_{j}$ is
@@
where $\lambda\asymp\sqrt{\log(q+p)/n}$ and $\mu\asymp\log n$ are two tuning parameters. Details on solving this constrained optimization and selecting the tuning parameters can be found in Section 4.1 of [23], where the constrained minimization problem is equivalently transformed to an unconstrained minimization problem with a readily available optimizer. Consequently, with $\hat{U}=(\hat{u}_{1},\ldots,\hat{u}_{q})\in\mathbb{R}^{(p+q)\times q}$, the estimated bias term in (7) is $n^{-1}\hat{U}^{\top}X^{\top}(X\hat{\theta}-Y)$, based on which we propose the following debiased estimator for $\gamma$,
@@
In the above, the constraint (9) is introduced to tackle the first term on the right-hand side of (8); this term may be viewed as the further bias associated with the estimator $n^{-1}u^{\top}X^{\top}(X\hat{\theta}-Y)$ for the $j$th coordinate $\tilde{g}_{j}^{\top}(\hat{\theta}_{M}-\theta_{M})$ of the bias in (7). This idea of minimizing the variance of a projected sum of errors while constraining the bias is a common practice in high-dimensional statistical inference with bias correction [18, 24]. [22] further introduced an additional constraint (10) to ensure that the variance of the second term on the right-hand side of (8) asymptotically dominates the first term for a fixed counterpart of $g_{j}$, and hence the variance of the debiased estimator would asymptotically dominate its bias. In our scenario, each $g_{j}$ is random due to estimating the parameter $\beta_{A}$ of the additional equation (4). This distinction, combined with the presence of two coupled equations in our model, sets our theoretical analysis apart from previous works. The constraint (11) is primarily a technical requirement to ensure that the error terms in (8) satisfy the Lindeberg’s condition.